### v1

# eTraM: Event-based Traffic Monitoring Dataset

Abstract
Event cameras, with their high temporal and dynamic range and minimal memory usage, have found applications in various fields. However, their potential in static traffic monitoring remains largely unexplored. To facilitate this exploration, we present eTraM - a first-of-its-kind, fully event-based traffic monitoring dataset. eTraM offers of data from different traffic scenarios in various lighting and weather conditions, providing a comprehensive overview of real-world situations. Providing 2M bounding box annotations, it covers eight distinct classes of traffic participants, ranging from vehicles to pedestrians and micro-mobility. eTraM’s utility has been assessed using state-of-the-art methods for traffic participant detection, including RVT, RED, and YOLOv8. We quantitatively evaluate the ability of event-based models to generalize on nighttime and unseen scenes. Our findings substantiate the compelling potential of leveraging event cameras for traffic monitoring, opening new avenues for research and application. eTraM is available at https://eventbasedvision.github.io/eTraM.
1 Introduction
In the dynamic landscape of modern transportation, Intelligent Transportation Systems (ITS) play a pivotal role in enhancing traffic flow [42, 24], route optimization [22, 39], and safety [15, 12]. As an essential task in ITS, traffic participant detection aims to provide information assisting in counting, measuring speed, identifying traffic incidents, and predicting traffic flow. The detection methods must be fast enough to operate in real-time and be insensitive to illumination change and varying weather conditions while utilizing less storage. For instance, in the span of just , a vehicle may travel over , and a pedestrian could cover over , leading to potential misses in fast-paced traffic scenarios and introducing motion blur concerns [40]. Moreover, nighttime and different weather conditions make the detection task more challenging as many features such as edge, corner, and shadow do not work due to varying illumination. In the face of these challenges, the integration of event cameras into ITS holds great promise for robust traffic participant detection in real-time scenarios.
Event cameras capture an asynchronous and continuous stream of “events" or pixel-level brightness changes instead of traditional static frames at fixed frequencies. Each event is represented by a tuple corresponding to an illuminance change greater than a fixed threshold at pixel location and time , with the polarity indicating whether the illuminance increased or decreased. Event cameras with their exceptional temporal resolution (over ) and high dynamic range (above ) [31] have prompted explorations into visual perception, robotics [14, 41], and its various applications in ITS [32, 37, 14].
Existing multimodal traffic datasets captured from various sensors, including RGB cameras, LiDAR, and Radar, have been utilized for several tasks in the context of autonomous vehicles (AV) [17, 36, 8, 9]. However, a largely unexplored yet promising domain lies in the use of event cameras for detection in traffic monitoring. This serves as an inspiration for us to contribute a first-of-its-kind, fully event-based traffic monitoring dataset.
In this paper, we present eTraM a novel fully event-based traffic perception dataset curated using the state-of-the-art (SOTA) high-resolution Prophesee EVK4 HD event camera [1]. The dataset spans over of annotated data from a static perspective that facilitates comprehensive traffic monitoring. We consulted experts from the Institute of Automated Mobility (IAM) [3] and strategically mounted an event camera over selected sites (intersections, roadways, and local streets) to collect traffic data under diverse conditions. The data collection process was systematically conducted across various weather and lighting conditions spanning challenging scenarios such as high glare, overexposure, underexposure, nighttime, twilight, and rainy days. eTraM includes over bounding box annotations of traffic participants such as vehicles (cars, trucks, buses, trams), pedestrians, and various micro-mobility (bikes, bicycles, wheelchairs) as shown in Figure 1.
eTraM offers the perspective of a static event camera captured at various scenes, further enhancing its versatility and applicability in real-world scenarios. This approach ensures that eTraM captures not only the routine dynamics of traffic but the nuances and challenges presented by a broad spectrum of scenarios and participants as well. We tested the diversity and quality of the dataset through various experiments and evaluated the generalization of event-based methods on nighttime and unseen scenes. eTraM stands as a valuable resource, propelling research and innovation in the evolving field of event-based traffic perception in ITS. Our contributions can be summarized as follows:
1.
A first-of-its-kind fully event-based dataset from a static perspective encompassing a diverse variety of traffic scenarios (scenes, weather, and lighting conditions) and participants (vehicles, pedestrians, and micro-mobility).
2.
Establish baselines using state-of-the-art event-based methods on eTraM across various traffic monitoring scenes and lighting conditions.
3.
Quantitative evaluation of the generalization capabilities of event-based methods on nighttime and unseen scenarios.
The rest of this paper is structured as follows. Section 2 briefly summarizes the existing event-based dataset contributions. Section 3 presents eTraM dataset with details on the acquisition framework, preprocessing, and annotation. In Section 4, we provide baseline results and explore the generalization capabilities of event-based methods on nighttime and unseen scenarios. Section 5 highlights the advantages of utilizing event cameras and the difference between static and ego-motion event datasets.
2 Related Datasets
| Dataset Name | Year | Duration (in hours) | Perspective |
Lighting | Weather | No. of Bbox | Scenarios | ||||||||
| Ego | Static | VH | PED | MM | Day | Night | Twilight | Clear | Rainy | |||||||
| DDD17 [4] | 2017 | 12 | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | - | Driving | |||||
| MVSEC [43] | 2018 | - | ✓ | ✓ | ✓ | ✓ | - | Driving, Handheld | ||||||||
| DVS Pedestrian [25] | 2019 | 0.1 | ✓ | ✓ | ✓ | ✓ | ✓ | 4.6K | Walking street | |||||||
| DDD20 [18] | 2020 | 51 | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | - | Driving | ||||||
| Gen1 [11] | 2020 | 39 | ✓ | ✓ | ✓ | ✓ | ✓ | 255K | Driving | |||||||
| 1 Megapixel [27] | 2020 | 15 | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | 25M | Driving | |||||
| DSEC [16] | 2021 | 1 | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | 390K | Driving | |||||
| DVS-OUTLAB [5] | 2021 | 7 | ✓ | ✓ | ✓ | ✓ | ✓ | 47K | Playground | |||||||
| PEDRo [6] | 2023 | 0.5 | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | 43K | Robotics | |||||
| eTraM (Ours) | 2024 | 10 | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | 2M |
Early event-based datasets often involved the transformation of frame-based datasets into event streams. A noteworthy example is [26], where MNIST [20] and Caltech-101 [13] datasets were converted to event streams by moving an event camera in front of a screen displaying frames. Later works utilized event simulators [30, 19] to convert widely-used frame-based datasets into their event-based counterparts.
Ego-motion event-based datasets have seen a rise in recent years due to the increased accessibility of event sensors. DDD17 [4], DDD20 [18] pioneered the initial efforts in deploying event cameras for driving scenarios using a DAVIS sensor. These datasets, focused on steering angle prediction, have and of driving data, respectively. MVSEC [43] presents a multimodal stereo dataset using DAVIS sensors along with LiDARs, IMUs, and RGB cameras for three-dimensional perception tasks, such as feature tracking, visual odometry, and stereo depth estimation. DSEC [16] further expands these fusion efforts by including 390K annotations for detection tasks on an hour of multimodal stereo data using Prophesee Gen3.1 sensors.
Prophesee [28] introduced two substantial ego-motion datasets [11, 27] for detection. Gen1 Automotive Detection Dataset [11] encompasses a total of 255,781 manually annotated bounding boxes (228,123 cars and 27,658 pedestrians instances) acquired over a span of using Prophesee Gen1 sensor. 1 Megapixel Automotive Dataset [27] stands out as the most comprehensive ego-motion event-based detection dataset. It encompasses of recorded footage, featuring a resolution of , with 25M generated bounding boxes. However, they are unable to provide nighttime annotations due to their automated labeling protocol. PEDRo dataset [6] focuses on people detection from a robotics perspective containing 43,259 bounding boxes from 119 recordings with an average duration of .
Static perception event-based datasets such as DVS-Pedestrian [25] is limited to pedestrian detection. It has 12 sequences recorded using a DAVIS346 camera and contains 4670 labeled instances of pedestrians. DVS-OUTLAB [5] explores the plausibility of using event cameras for long-time monitoring. It consists of recordings from three static CeleX-4 DVS event cameras featuring outdoor urban public areas involving persons, dogs, bicycles, and sports balls as objects of interest.
Table 1 provides an overview of existing event-based datasets. Interest in event cameras has been piqued for intelligent transportation, with many datasets focused on the ego-vehicle perspective, but there is a lack of datasets focusing on traffic monitoring from a static perspective.
3 eTraM Dataset
This section details eTraM’s acquisition framework, preprocessing techniques, annotation, and statistics, providing insights into its applicability for traffic monitoring.
3.1 Dataset Acquisition Framework
We rely on the Prophesee EVK4 HD event camera [1], notable for its high resolution (), temporal resolution (over ), dynamic range (above ), and exceptional low light cutoff (), to capture high-quality data. The event camera was strategically positioned at approximately with a pitch angle of about to the ground. This configuration is deliberately chosen to maintain consistency with the placement of traffic cameras and to ensure comprehensive coverage of interactions between diverse traffic participants. Figure 2 shows different sites considered for data collection.
eTraM comprises recorded sequences from multiple intersections, roadways, and local streets around Arizona State University, Tempe campus. The sequences were recorded for intervals of at different times of the day, covering daytime, nighttime, and twilight. The dataset also observes different weather conditions, including sunny, overcast, and rainy. To achieve this, extensive data collection efforts were carried out over a span of .
3.2 Preprocessing and Annotation
Given the sensitive nature of the event sensor, it was observed that nighttime data tends to exhibit higher levels of noise, primarily attributed to reflections and pointed sources of light from streets and vehicles. To address this challenge and enhance the quality of our data, the recorded sequences are passed through a spatiotemporal filter [7]. This spatiotemporal filter works on the idea that events from real objects should occur closer together in both space and time more often compared to events from random noises [14].
The filtering mechanism discards an event if a threshold amount of events with the same polarity do not occur within a fixed temporal window in the vicinity of its 8-neighborhood spatial coordinates. For any , this condition is represented mathematically in Equation 1,
where t represents the temporal window and P is 1 only when the polarity of and are the identical. We choose a temporal window of with a minimum threshold of 2 neighboring events for our experiments. The specific filter values were determined through experiments detailed in [5] and further validated through a smaller experiment conducted during the training phase. We illustrate the significance of spatiotemporal filtering through visualization in the supplementary material.
Following the denoising stage, events within the stream are partitioned into discrete time bins and consolidated into a single frame, thereby converting the asynchronous event stream into synchronous frames of . These frames are then annotated using CVAT [33], an open-source annotation tool. Our rigorous manual annotation process resulted in the precise identification of over 2M 2D bounding boxes.
| Traffic Site | Lighting | RVT | RED | YOLOv8 | |||||||||
| PED | VH | MM | All | PED | VH | MM | All | PED | VH | MM | All | ||
| Intersections | Daytime | 0.460 | 0.813 | 0.315 | 0.722 | 0.395 | 0.593 | 0.284 | 0.545 | 0.167 | 0.293 | 0.111 | 0.190 |
| Roadways | 0.430 | 0.733 | 0.070 | 0.627 | 0.347 | 0.590 | 0.055 | 0.551 | 0.173 | 0.290 | 0.004 | 0.156 | |
| Local Streets | 0.196 | 0.938 | 0.586 | 0.316 | 0.208 | 0.875 | 0.695 | 0.351 | 0.124 | 0.559 | 0.204 | 0.296 | |
| All Scenes | 0.304 | 0.781 | 0.403 | 0.572 | 0.302 | 0.656 | 0.251 | 0.497 | 0.142 | 0.309 | 0.112 | 0.188 | |
| Intersections | Nighttime | 0.161 | 0.465 | - | 0.262 | 0.149 | 0.425 | - | 0.242 | 0.071 | 0.375 | - | 0.149 |
| Roadways | 0.310 | 0.827 | - | 0.739 | 0.362 | 0.782 | - | 0.726 | 0.004 | 0.229 | - | 0.117 | |
| Local Streets | 0.739 | 0.868 | 0.097 | 0.829 | 0.722 | 0.831 | 0.145 | 0.817 | 0.198 | 0.486 | 0.030 | 0.239 | |
| All Scenes | 0.317 | 0.674 | 0.064 | 0.523 | 0.303 | 0.660 | 0.083 | 0.504 | 0.123 | 0.322 | 0.013 | 0.153 | |
| Overall | 0.309 | 0.717 | 0.313 | 0.539 | 0.303 | 0.649 | 0.197 | 0.491 | 0.134 | 0.314 | 0.086 | 0.178 | |
3.3 Dataset Statistics
Here, we highlight key characteristics of the collected data and annotations. The dataset encompasses three distinct traffic monitoring scenes with of intersection, of roadway, and of local street data sequences. Data for each scene is collected at multiple locations. For instance, the intersection scene contains data from 2 four-way, three-way, and an uncontrolled intersection. Each location has daytime, nighttime, and twilight data totaling up to of data with of daytime and nighttime data.
When comparing the event distribution in eTraM with other ego-motion event-based traffic datasets like 1 Megapixel Automotive and DSEC for a duration of , we observe that the number of events in eTraM is significantly lesser by a factor of 30, as shown in Figure 3 (a). This is accredited to the event camera being static in eTraM, primarily focusing on the dynamic traffic participants in a scene. In contrast, datasets from an ego-motion perspective capture more data due to the surrounding infrastructure’s relative motion and lead to dense streams of events. The sparsity of events in eTraM, combined with the asynchronous nature of events, leads to low memory utilization, which is particularly advantageous for traffic monitoring infrastructure.
eTraM contains over 2M instances of 2D bounding box annotations for traffic participant detection. These annotations additionally include object IDs, making it possible to evaluate multi-object tracking, as shown in the supplementary material. The annotation classes encompass a range of traffic participants, from various vehicles (cars, trucks, buses, and trams) to pedestrians and micro-mobility (bikes, bicycles, and wheelchairs).
Figure 3 (b) illustrates the spatial distribution of each class within the frame. We observe a uniform distribution of vehicle instances across the entire frame, while instances belonging to the pedestrian class cover more than 50% of the frame. This safeguards the model from developing a bias for classes in a specific region of the frame.
For accessibility and ease of use, eTraM is provided in multiple formats: RAW, DAT, and H5 [29]. Additionally, annotations are available in numpy format. The dataset is split into 70% training, 15% validation, and 15% testing, ensuring each subset has proportional data from each scene. Further statistics of eTraM can be seen in Figure 3 (c) & 3 (d). To the best of our knowledge, this stands as a first-of-its-kind event-based dataset for traffic monitoring. Additionally, the inclusion of nighttime data enhances its versatility for a broader range of research applications.
4 Dataset Evaluation
In this section, we present the results of experiments evaluating eTraM for real-world detection scenarios. We first train and evaluate models to establish baselines and assess their performance in various scenarios (Section 4.1). We further conduct experiments to test the ability of these models to generalize on nighttime conditions and unseen scenes (Section 4.2). For all experiments, we evaluate eTraM on SOTA tensor-based object detection methods, specifically Recurrent Vision Transformers (RVT), Recurrent Event-camera Detector (RED), and a frame-based method, You Only Look Once (YOLOv8). This comparison helps us understand how SOTA utilizing dense tensor representation, performs compared to the conventional frame-based method. For evaluation, we focus on the three broad categories of traffic participants, vehicles (VH), pedestrians (PED), and micro-mobility (MM), and report the mean Average Precision at an intersection-over-union (IoU) threshold of 50% (AP50).
4.1 Baseline Evaluation
To assess the performance of detectors of eTraM, we train them on of data and evaluate them on of validation and of test data. RVT and RED were trained from scratch over on A100, while YOLOv8 reached convergence in . We conducted separate evaluations to provide insights into how each model performs in different scenes and lighting conditions. This allows us to understand how these models handle diverse and changing environmental contexts.
Several key observations emerged from the evaluation results of tensor-based models shown in Table 2. Notably, the performance of vehicle detection consistently outperforms that of pedestrians across all scenes and models. The performance in the micro-mobility category exhibits significant variance, likely due to the smaller number of instances and the broad range of subjects captured within this category. During daytime, both vehicle and pedestrian detection results are consistent in intersections and roadways. However, in local streets, the performance of vehicle detection improves compared to other scenes due to fewer instances and reduced occlusion. Conversely, we observe a decline in the performance of pedestrian detection due to higher pedestrian densities leading to more inter-class occlusions.
During nighttime, the performance of vehicle detection in local streets and roadways remains consistent. Interestingly, pedestrian detection results improve significantly on local streets at night. This can be attributed to factors such as reduced pedestrian-to-pedestrian occlusion, instances being closer to the camera, and additional visual features like shadows that become more prominent at night due to street lights. Due to noise from various light sources, a drop in performance is observed at nighttime intersections. Despite this, the performance during nighttime is at par with daytime scenarios, with an increase in performance on vehicle detection on roadways at night. This increase could be due to reduced inter-class occlusion as we observe fewer vehicles in nighttime conditions.
Furthermore, tensor-based methods such as RED and RVT, which incorporate temporal information through Recurrent Neural Networks in their architecture, consistently outperform the conventional frame-based method of YOLOv8. Figure 4 shows an example of detection on eTraM. The evaluation highlights the challenges and strengths of various traffic monitoring scenarios and categories, particularly showcasing the effectiveness of event-based models in nighttime conditions.
4.2 Generalization Evaluation
A key characteristic required for real-world deployment is that models demonstrate transferability to unseen scenarios. Since event cameras are invariant to absolute illuminance levels, we expect event-based models to also demonstrate transferability to nighttime data. We conduct experiments where we control the train and test set to evaluate their transferability in the following sections quantitatively.
4.2.1 Generalization on Night time
In [27], qualitative assessments were conducted to compare the generalization capabilities of event-based models against frame-based models in nighttime scenarios. We aim to quantitatively assess how well an event-based model trained on daytime data can perform in nighttime conditions. We conducted a controlled experiment where we trained two models for each architecture, one on of only daytime data and another with of daytime along with of nighttime data. The evaluation of both models was conducted on previously unseen nighttime data.
| Train Set | RVT | RED | ||
| VH | PED | VH | PED | |
| Day | 0.566 | 0.166 | 0.374 | 0.354 |
| Day+Night | 0.761 | 0.254 | 0.673 | 0.422 |
The summarized results can be found in Table 3. We observe a consistent trend across every object class and architecture: models trained on data with nighttime sequences consistently outperform models trained solely on daytime sequences. Despite the expectation that event cameras would exhibit proficient performance in nighttime scenarios, these observations reveal that relying solely on daytime event data may not be sufficient. These models fail to attain performance levels comparable to those achieved by models trained on nighttime data with the model. This suggests that the unique challenges posed by nighttime conditions necessitate explicit training with relevant data to ensure optimal model performance. Upon closer inspection, we find that nighttime data has environmental interferences and distinct variations of noise. This could be inherent in nighttime data due to the heightened sensitivity of event cameras, which is a factor in this discrepancy.
While event-based models are unable to generalize on nighttime data out of the box, they have at par results on daytime and nighttime data when trained on the combined dataset (refer Table 2). This highlights the need for labeled nighttime data to allow event-based models to augment current frame-based systems.
4.2.2 Generalization on Unseen Scenes
To substantiate the ability of event-based detectors to generalize to previously unencountered traffic scenes, we train our model on a subset of the dataset. We evaluate the ability by testing each architecture on two independent test sets, one “held in" that contains sequences from intersections that our model has seen during the training phase and the other on the “held out" test set with data from an intersection that is skipped during training.
| Test Set | RVT | RED | ||
| VH | PED | VH | PED | |
| Held In | 0.449 | 0.316 | 0.556 | 0.521 |
| Held Out | 0.628 | 0.529 | 0.572 | 0.509 |
As depicted in Table 4, we observe a comparable performance of the models across both the major categories considered. On evaluating the model on the “held out" test set, representing an entirely new and unseen traffic scene, the model’s generalization capability is evident. These values are at par with the performance on the “held in" test set. This similarity in results highlights the model’s ability to seamlessly extend its learning to previously unseen traffic scenes, validating its transfer capability across changing environments. This generalization to unseen intersections is pivotal for real-world deployment.
5 Discussion
5.1 Event Cameras in Traffic Monitoring
In this section, we discuss the advantages of using event cameras to augment traditional frame camera systems for traffic monitoring.
Traditional cameras take continuous snapshots at a fixed frequency, potentially capturing individuals’ identities and sensitive information. In comparison, event cameras register events with no color and texture information, significantly reducing the probability of gathering sensitive information of any individual [27, 5]. Further, the advantages of event cameras extend to their performance in low light conditions and their robustness towards mitigating motion blur. As we demonstrate through our baseline evaluations in Table 2, event cameras display equally superior performance in nighttime conditions while maintaining a high temporal resolution that aids it in substantially reducing motion blur [31]. An essential requirement of traffic monitoring is the need for resource-efficient sensors. Since event cameras are designed to operate on a sparse data stream generated by significant visual changes, they exhibit lower memory requirements and power consumption compared to traditional frame cameras [2, 34]. This makes event cameras sustainable and cost-effective for continuous monitoring over extended periods.
We further highlight the capabilities brought by event cameras in traffic monitoring scenarios through two practical applications, as illustrated in Figure 5. Firstly, in detecting stop sign violations, event cameras excel in capturing instantaneous changes in the visual scene, enabling precise detection of vehicles coming to a halt. The number of events associated with the object makes it easy to discern whether a moving vehicle is slowing down or has stopped. Secondly, due to the high temporal resolution of the event camera and the event density corresponding to an object, it is straightforward to detect sudden acceleration or erratic behavior of fast-moving traffic participants. This capability is critical for identifying potentially hazardous situations on the road, such as near misses, and allowing for prompt intervention or alert generation.
5.2 Static and Ego Event-based Datasets
Next, we discuss the difference between static and ego-motion event-based datasets, emphasizing the significance of the former for traffic monitoring. In static datasets events are more likely to be associated with relevant traffic participants rather than extraneous background changes seen in ego-motion data, enhancing the visibility of these participants. This enhanced visibility is illustrated in Figure 6, which provides a qualitative comparison between events captured from a static and an ego-motion event camera. As shown, dynamic traffic participants from the static camera are better visible than those from the ego-motion camera. Another distinction is the field of view provided by static roadside datasets. Positioned at an elevated height, these cameras capture a broader scene, including far-away objects, while ego-motion datasets are constrained to the vehicle’s immediate surroundings. The elevated perspective of static datasets proves advantageous, enabling the detection of traffic participants at a distance and providing a comprehensive view of the traffic environment, as seen in Figure 7. This would be beneficial to a broad spectrum of applications, like traffic scene estimation [23, 38, 21], road safety monitoring [35] and traffic signal control [10].
6 Conclusion
We present eTraM, a large-scale manually annotated event-based dataset for traffic monitoring. Our meticulously curated dataset, captured using the cutting-edge Prophesee EVK4 HD high-resolution event camera, provides new traffic detection and tracking opportunities from a static perspective. With over of annotated event data spanning various scenes and lighting conditions, our dataset offers insights into the complex dynamics of traffic scenarios. The comprehensive annotations include over 2M bounding boxes of various traffic participants, from vehicles to pedestrians and micro-mobility. Further, nighttime generalization results highlight the value of labeled nighttime data, enabling event-based models to achieve daytime-equivalent performance on nighttime data. We demonstrate the ability of event-based models to generalize effectively to unseen scenarios, emphasizing its potential in real-world traffic monitoring. As the field of event-based sensing continues to advance, eTraM holds the potential to serve as an invaluable resource in driving its development and presenting new opportunities in intelligent transportation systems.
7 Acknowledgment
This research is sponsored by NSF Pathways to Enable Open-Source Ecosystems grant (#2303748), the Partnerships for Innovation grant (#2329780), the CPS grant (#2038666), CRII grant (# 2153311). We thank data collection support from the Institute of Automated Mobility.

### v2

@@
-eTraM: Event-based Traffic Monitoring Dataset
+scale=1, color=gray, angle=0, opacity=0.5, position=current page.north, vshift=-1cm, contents= This paper has been accepted for publication at the IEEE Conference on Computer Vision and Pattern Recognition (CVPR), Seattle, 2024.
+©IEEE eTraM: Event-based Traffic Monitoring Dataset
 Abstract
@@
 4.1 Baseline Evaluation
-To assess the performance of detectors of eTraM, we train them on of data and evaluate them on of validation and of test data.
-RVT and RED were trained from scratch over on A100, while YOLOv8 reached convergence in .
+To assess the performance of detection methods on eTraM, we train them on of data and evaluate them on of validation and of test data.
+RVT and RED were trained from scratch over on A100, while YOLOv8 was trained for .
 We conducted separate evaluations to provide insights into how each model performs in different scenes and lighting conditions.
@@
 The performance in the micro-mobility category exhibits significant variance, likely due to the smaller number of instances and the broad range of subjects captured within this category.
-During daytime, both vehicle and pedestrian detection results are consistent in intersections and roadways.
-However, in local streets, the performance of vehicle detection improves compared to other scenes due to fewer instances and reduced occlusion.
+During the daytime, both vehicle and pedestrian detection results are at par in intersections and roadways.
+In local streets, the performance of vehicle detection improves compared to other scenes due to fewer instances and reduced occlusion.
 Conversely, we observe a decline in the performance of pedestrian detection due to higher pedestrian densities leading to more inter-class occlusions.
@@
 This can be attributed to factors such as reduced pedestrian-to-pedestrian occlusion, instances being closer to the camera, and additional visual features like shadows that become more prominent at night due to street lights.
-Due to noise from various light sources, a drop in performance is observed at nighttime intersections.
+Due to noise from various light sources, a drop in performance is observed at intersections during nighttime.
 Despite this, the performance during nighttime is at par with daytime scenarios, with an increase in performance on vehicle detection on roadways at night.
@@
 Figure 4 shows an example of detection on eTraM.
-The evaluation highlights the challenges and strengths of various traffic monitoring scenarios and categories, particularly showcasing the effectiveness of event-based models in nighttime conditions.
+The baseline results highlight the challenges and strengths of various traffic monitoring scenarios and categories, particularly showcasing the effectiveness of event-based models in nighttime conditions.
 4.2 Generalization Evaluation
@@
 In [28], qualitative assessments were conducted to compare the generalization capabilities of event-based models against frame-based models in nighttime scenarios.
-We aim to quantitatively assess how well an event-based model trained on daytime data can perform in nighttime conditions.
+We aim to quantitatively assess how well event-based models RVT and RED, trained on daytime data, can perform in nighttime conditions.
 We conducted a controlled experiment where we trained two models for each architecture, one on of only daytime data and another with of daytime along with of nighttime data.
-The evaluation of both models was conducted on previously unseen nighttime data.
+Both models were evaluated on previously unseen nighttime data.
 | Train Set | RVT | RED | ||
@@
 The summarized results can be found in Table 3.
-We observe a consistent trend across every object class and architecture: models trained on data with nighttime sequences consistently outperform models trained solely on daytime sequences.
+We observe a consistent trend across every object class and architecture where models trained on data with nighttime sequences consistently outperform models trained solely on daytime sequences.
 Despite the expectation that event cameras would exhibit proficient performance in nighttime scenarios, these observations reveal that relying solely on daytime event data may not be sufficient.
@@
 | Held Out | 0.628 | 0.529 | 0.572 | 0.509 |
-As depicted in Table 4, we observe a comparable performance of the models across both the major categories considered.
+As depicted in Table 4, we observe a comparable performance of the models across both major categories.
 On evaluating the model on the “held out" test set, representing an entirely new and unseen traffic scene, the model’s generalization capability is evident.
 These values are at par with the performance on the “held in" test set.
-This similarity in results highlights the model’s ability to seamlessly extend its learning to previously unseen traffic scenes, validating its transfer capability across changing environments.
+This similarity in results highlights the model’s ability to seamlessly extend its learning to previously unseen traffic scenes, validating its transferability across changing environments.
 This generalization to unseen intersections is pivotal for real-world deployment.
@@
 Traditional cameras take continuous snapshots at a fixed frequency, potentially capturing individuals’ identities and sensitive information.
-In comparison, event cameras register events with no color and texture information, significantly reducing the probability of gathering sensitive information of any individual [27, 5].
+In comparison, event cameras register events with no color and texture information, significantly reducing the probability of gathering sensitive information of any individual [28, 5, 9].
 Further, the advantages of event cameras extend to their performance in low light conditions and their robustness towards mitigating motion blur.
@@
 Another distinction is the field of view provided by static roadside datasets.
-Positioned at an elevated height, these cameras capture a broader scene, including far-away objects, while ego-motion datasets are constrained to the vehicle’s immediate surroundings.
+Positioned at an elevated height, these cameras capture a broader scene and include far-away objects, while ego-motion datasets are constrained to the vehicle’s immediate surroundings.
 The elevated perspective of static datasets proves advantageous, enabling the detection of traffic participants at a distance and providing a comprehensive view of the traffic environment, as seen in Figure 7.